Read and analyze the given problem. Provide a complete solution, Table, and a graph corresponds to the answer.



According to last year’s report, a filipino household spends an average of P400/day for food. Suppose you took a sample of 25 households. You determined how much each household spent for the food daily and the results revealed a mean of P380 and a standard deviation of P21.50. with 99% confidence what would be your conclusion




1
Expert's answer
2021-06-04T10:43:10-0400

One sample t-test.

Null hypothesis "H_0:\mu=400."

Alternative hypothesis "H_a:\mu\ne400."

Degrees of freedom: "df=n-1=25-1=24."

Test statistic: "t=\frac{\bar x-\mu}{\frac{s}{\sqrt{n}}}=\frac{380-400}{\frac{21.50}{\sqrt{25}}}=-4.65."

P-value: "p=2P(T<-4.65)=0.0001."

Since the P-value is less than 0.01, reject the null hypothesis.

There is a significant evidence that the average Filipino household spend on foot is different from P400.